# Holonomy of the Obata connection on 2-step hypercomplex nilmanifolds

### v1

@@ 6 Holonomy of $2$-step hypercomplex nilmanifolds
This provides an explicit example of a non-trivial, abelian holonomy algebra arising from a non-flat Obata connection on a $2$-step nilmanifold.
@@
Let $G$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. If the Obata connection is flat, then the full holonomy group is discrete. A natural question is whether this could be trivial. To address this question, we prove the following theorem.
@@
Let $(\Gamma\backslash G,\{J_{\alpha}\})$ be a $2$-step nilmanifold endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$ such that the Obata connection is flat. Then the holonomy group $\operatorname{Hol}(\nabla)$ is trivial if and only if $\G\Gamma\backslash$ G is a torus.
  • Proof.
@@
    The Obata connection is complex with respect to each $J_{\alpha}$ and has vanishing (and hence parallel) torsion. Moreover, as shown in the proof of Theorem 6.1, for any $X\in\mathfrak{g}$, the endomorphism $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$ is nilpotent (see Equation (6.2)). Since the Obata connection $\nabla$ is flat and torsion-free, it follows from [Kim, Theorem 2.2 and Proposition 1.2] that $\nabla$ is complete.
    We apply [ABD, Theorem 4.1] to conclude that $(\Gamma\backslash G,J_{\alpha},\nabla)$ is affinely biholomorphic to $(\Gamma_{0}\backslash G_{0},J_{\alpha}^{0},\nabla^{0})$ for each $\alpha$, where $G_{0}$ is the connected and simply connected Lie group whose Lie algebra is given by
[…]
    $\Gamma_{0}$ is a lattice in $G_{0}$, $J_{\alpha}^{0}$ is the complex structure induced on $P^{\nabla}$ by $J_{\alpha}$ and $\nabla^{0}$ is the ()(-)-connection.
[…]
@@
Theorem 6.7 still holds true if instead of considering a 2-step nilmanifold we assume that $\G\Gamma\backslash$ G is a solvmanifold with $G$ a completely solvable almost abelian Lie group. Indeed, in this case the Obata connection is automatically flat, according to [AB1, Proposition 3.7], and is complete, due to [AB1, Corollary 3.8]. For the last paragraph, we use the Saito rigidity Theorem [Sai], which extends the result of Mal’cev from nilpotent to completely solvable Lie groups.
@@ 7 Finale
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a Lie algebra endowed with a hypercomplex structure such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for all $\alpha=1,2,3$. We claim that $\mathfrak{g}$ is at most $3$-step nilpotent. Indeed, using the integrability of $J_{\alpha}$ and the assumption $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, one obtains
$X,Y],Z]=-J_{\alpha}[[X,Y],J_{\alpha}Z],$ (7.1)
@@
Since in the $2$-step case the corresponding Obata connection is flat (see Corollary 3.2), we will focus our attention on the $3$-step case. We observe that the hypercomplex structure must be non-abelian. Indeed, if $\{J_{\alpha}\}$ is abelian and $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, then
@@
Examples of 3-step nilpotent Lie algebras with hypercomplex structures have been given in [DF3, BD, AB2]. Among these, the only examples with non-abelian hypercomplex structures are those in [AB2]; however, those are Obata flat and none of them satisfies the condition $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for all $\alpha=1,2,3$.
Due to the lack of examples of this kind in the literature, we include here a 3-step nilpotent example satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for any $\alpha=1,2,3$.
@@
$\begin{split}&de^{i}=0,\ i=1,\dots,4,\ \ de^{5}=e^{12}-e^{34},\\ &de^{i}=0,\ i=6,7,8,\ \ de^{9}=e^{13}-e^{42},\\ &de^{i}=0,\ i=10,11,12,\ \ de^{13}=-e^{25}+e^{39},\\ &de^{14}=e^{15}+e^{49},\ \ de^{15}=e^{45}-e^{19},\ \ de^{16}=-e^{35}-e^{29}.\end{split}$
@@
satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, for each $\alpha=1,2,3$.
Proposition 7.1.
Let $(\mathfrak{g},\{J_{\alpha}\})$ be a hypercomplex Lie algebra such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha=1,2,3$. Then
$R(X,Y)=-\operatorname{ad}_{[X,Y]},$
and $(\mathfrak{g},\{J_{\alpha}\})$ is non-flat if and only if $\mathfrak{g}$ is $3$-step nilpotent.
Let $G$ be the corresponding connected and simply connected $3$-step nilpotent Lie group. Then the holonomy algebra of the Obata connection on $G$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$.
  • Proof.
    As we pointed out at the beginning of this section, $\mathfrak{g}$ is at most $3$-step nilpotent. Applying Formula (2.2) and the hypothesis that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha$, we get that
    $\nabla_{X}\nabla_{Y}Z$ $=\frac{1}{4}\left([X,[Y,Z]+J_{1}[J_{1}X,[Y,Z]]-J_{2}[X,[Y,J_{2}Z]]+J_{3}[J_{1}X,[Y,J_{2}Z]]\right)$
    $=0,$
    where we used Equation (7.1). Therefore,
    $R(X,Y)Z=-\nabla_{[X,Y]}Z=-\frac{1}{2}([[X,Y],Z]-J_{2}[[X,Y],J_{2}Z])=-[[X,Y],Z],$
    where the last equality again follows from Equation (7.1).
    Set $\mathfrak{a}:=\text{span}\{\operatorname{ad}_{[X,Y]}\ |\ X,Y\in\mathfrak{g}\}$. Then, $\mathfrak{a}$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$. In fact,
    $\operatorname{ad}_{[X,Y]}\operatorname{ad}_{[Z,W]}U=[[X,Y],[[Z,W],U]]=0,$
    as $\mathfrak{g}$ is $3$-step nilpotent.
    Furthermore, using that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, $\mathfrak{g}^{2}=[\mathfrak{g},\mathfrak{g}^{1}]\subset\mathfrak{z}$ and $\mathfrak{g}^{1},\mathfrak{g}^{1}]=0$,
    $2\nabla_{U}\operatorname{ad}_{[X,Y]}Z$ $=[U,[[X,Y],Z]]+[J_{1}U,[[X,Y],Z]]$
    $\ \ \ \ -J_{2}[U,J_{2}[[X,Y],Z]]+J_{3}[J_{1}U,J_{2}[[X,Y],Z]]=0,$
    and
    $2\operatorname{ad}_{[X,Y]}\nabla_{U}Z$ $=[[X,Y],[U,Z]]=0.$
    By Theorem 6.5, $\mathfrak{a}=\mathfrak{hol}(\nabla)$, concluding the proof. ∎

### v2

@@ 6 Holonomy of 2-step hypercomplex nilmanifolds
This provides an explicit example of a holonomy algebra arising from a non-flat Obata connection on a $2$-step nilmanifold such that $\mathfrak{hol}(\nabla)$ is abelian and non-trivial.
The next lemma is one of the key ingredients in the proof of Theorem 6.9 below.
Lemma 6.5.
Let $G^{4n}$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. Then for any $X,Y\in\mathfrak{g}$, $\nabla_{X}Y\in\mathfrak{g}_{1}^{\mathbb{H}}$ and $\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}}\subset\mathfrak{h}$. In particular, for any $\in\mathfrak{g}, X\nabla_{X}$ is nilpotent.
The first claim is straightforward by Formula (2.2).
For the second claim, since $\nabla_{X}J_{\alpha}=J_{\alpha}\nabla_{X}$ and $\mathfrak{h}$ is $\{J_{\alpha}\}$-invariant, it suffices to prove that $\nabla_{X}\mathfrak{g}^{1}\subset\mathfrak{h}$. Applying Formula (2.2), we see that for any $Y\in\mathfrak{g}^{1}$:
$\nabla_{X}Y=\frac{1}{2}\big(-J_{2}[X,J_{2}Y]+J_{3}[J_{1}X,J_{2}Y]\big).$
With the same argument used in the proof of Step 1 in Theorem 6.1, the first term lies in $\mathfrak{g}^{1}\cap J_{2}\mathfrak{g}^{1}$, while the second term lies in $J_{1}\mathfrak{g}^{1}\cap J_{3}\mathfrak{g}^{1}$. The proof follows by the definition of $\mathfrak{h}$.
The fact that $\nabla_{X}$ is nilpotent follows since $(\nabla_{X})^{3}=0.$ In fact for any $Y\in\mathfrak{g}$
$\nabla_{X}(\nabla_{X}(\nabla_{X}Y))\in\nabla_{X}(\nabla_{X}\mathfrak{g}_{1}^{\mathbb{H}})\subset\nabla_{X}\mathfrak{h}=0,$
by Step 2 in the proof of Theorem 6.1. ∎
6.1 Trivial Obata holonomy
@@
Let $G$ be a $2$-step nilpotent Lie group endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$. If the Obata connection is flat, then the full holonomy group is discrete. A natural question is when the holonomy group is trivial; in this section we prove in Theorem 6.9 below that it is never the case.
In order to prove this, we recall some useful facts on affine connections. Let $M$ be a connected $n$-dimensional manifold with an affine connection $\nabla$. We define the following subspace of the space $\mathfrak{X}(M)$ of smooth vector fields on $M$:
$\mathcal{P}^{\nabla}=\{X\in\mathfrak{X}(M)\mid\nabla X=0\}.$ (6.2)
The subspace $\mathcal{P}^{\nabla}$ is in general neither finite dimensional as an $\mathbb{R}$-vector space nor closed under the Lie bracket of vector fields. However, under additional assumptions both conditions are satisfied, as the following well known lemma shows (see [Wo, Proposition 2.2] for the first item and [Wo, p. 323] for the second one).
Lemma 6.6 ([Wo]).
Let $M$ be a connected $n$-dimensional manifold with an affine connection $\nabla$.
  1. $(1)$
    If the holonomy group $\operatorname{Hol}(\nabla)$ is trivial, then $\mathcal{P}^{\nabla}$ is an $n$-dimensional real vector space.
  2. $(2)$
    The space $\mathcal{P}^{\nabla}$ is a Lie subalgebra of $\mathfrak{X}(M)$ if and only if the torsion $T$ of $\nabla$ is parallel.
In case the manifold admits an almost complex structure $J$ and an affine connection $\nabla$ with trivial holonomy, we will make use of the following result.
Lemma 6.7 ([ABD, Lemma 3.2]).
Let $M$ be a connected manifold with an almost complex structure $J$. Assume that there exists an affine connection $\nabla$ on M such that $\operatorname{Hol}(\nabla)$ is trivial. Then the following conditions are equivalent:
  1. $(1)$
    $\nabla J=0$;
  2. $(2)$
    the space $\mathcal{P}^{\nabla}$ of parallel vector fields on $M$ is $J$-invariant, that is, $J(\mathcal{P}^{\nabla})=\mathcal{P}^{\nabla}$.
Any Lie group $H$ admits a natural left-invariant connection with trivial holonomy: it is defined by $\nabla^{0}_{X}Y=0$ for all $X,Y$ left-invariant vector fields on $H$, and it is known as the ()(-)-connection (see, for instance, [KN, Chapter X, Proposition 2.12]). Note that its torsion is given by $T^{0}(X,Y)=-[X,Y]$ for $X,Y$ left-invariant vector fields on $H$. Moreover, $\nabla^{0}$ is flat, complete, and clearly $\nabla^{0}T^{0}=0$ by definition of $\nabla^{0}$.
If $\Lambda\subset H$ is any discrete subgroup of $H$, then $\nabla^{0}$ on $H$ induces a unique connection, still denoted $\nabla^{0}$, on $\H\Lambda\backslash$ H such that the projection $\H\pi:H\to\Lambda\backslash$ H is an affine local diffeomorphism. This connection $\nabla^{0}$ on $\H\Lambda\backslash$ H is geodesically complete, has trivial holonomy and its torsion is parallel.
If $J_{0}$ is a left-invariant complex structure on $H$ then, by definition of $\nabla^{0}$, it follows that $\nabla^{0}J_{0}=0$. The complex structure $J_{0}$ on $H$ induces a unique complex structure, still denoted $J_{0}$, on $\H\Lambda\backslash$ H such that $\pi$ is a local biholomorphism. Since $\pi$ is also affine, we have that $\nabla^{0}J_{0}=0$ on $\H\Lambda\backslash$ H. Therefore, $(\Lambda\backslash H,J_{0})$ carries a geodesically complete connection $\nabla^{0}$ with trivial holonomy and parallel torsion such that $\nabla^{0}J_{0}=0$. See [ABD] for further details.
Theorem 6.8 ([ABD, Theorem 4.1]).
The triple $(M,J,\nabla)$ where $M$ is a connected manifold endowed with a complex structure $J$ and a connection $\nabla$ with $\nabla J=0$ and trivial holonomy is affinely biholomorphic to a triple $(\Lambda\backslash H,J_{0},\nabla^{0})$ as above if and only if $\nabla$ is geodesically complete and its torsion is parallel.
The Lie group $H$ is the connected and simply connected Lie group with Lie algebra $\mathcal{P}^{\nabla}$ defined in (6.2), and $J_{0}$ is the complex structure on $\H\Lambda\backslash$ H induced by the restriction of $J$ to $\mathcal{P}^{\nabla}$.
As a consequence of Lemmas 6.5, 6.6 and 6.7 and Theorem 6.8 we obtain the main result of this section.
@@ 2. $(2)$
Let $(\Gamma\backslash G,\{J_{\alpha}\})$ be an at most 2-step nilmanifold endowed with an invariant hypercomplex structure $\{J_{\alpha}\}$ such that the Obata connection $\nabla$ is flat. Then the holonomy group $\operatorname{Hol}(\nabla)$ is trivial if and only if $\G\Gamma\backslash$ G is a torus.
In particular, if $(\Gamma\backslash G,\{J_{\alpha}\})$ is a $2$-step nilmanifold endowed with an invariant hypercomplex structure such that the Obata connection $\nabla$ is flat, then the holonomy group $\operatorname{Hol}(\nabla)$ is never trivial.
@@
The Obata connection $\nabla$ satisfies $\nabla J_{\alpha}=0$ for all $\alpha$ and has vanishing (and hence parallel) torsion. By assumption, $\operatorname{Hol}(\nabla)$ is trivial, therefore, Lemma 6.6 implies that the space $\mathcal{P}^{\nabla}$ of smooth parallel vector fields on $\G\Gamma\backslash$ G is a $4n$-dimensional Lie algebra, with $4n:=\dim\Gamma\backslash G=\dim G$, and it follows from Lemma 6.7 that $J_{\alpha}(\mathcal{P}^{\nabla})=\mathcal{P}^{\nabla}$ for all $\alpha$. We denote by $H$ the connected and simply connected Lie group with Lie algebra $\mathcal{P}^{\nabla}$.
On the other hand, it follows from Lemma 6.5 that, for any $X\in\mathfrak{g}$, the endomorphism $\nabla_{X}:\mathfrak{g}\to\mathfrak{g}$ is nilpotent. As the Obata connection $\nabla$ is flat and torsion-free, it follows from [Kim, Theorem 2.2 and Proposition 1.2] that $\nabla$ is geodesically complete on $G$, hence on $\G\Gamma\backslash$ G.
Since $\nabla$ is flat, torsion-free, geodesically complete and $\nabla J_{\alpha}=0$ for all $\alpha$, we may apply Theorem 6.8 to conclude that $(\Gamma\backslash G,J_{\alpha},\nabla)$ is affinely biholomorphic to $(\Lambda\backslash H,J_{\alpha}^{0},\nabla^{0})$ for each $\alpha$, where $\Lambda$ is a lattice in $H$, $J_{\alpha}^{0}$ is the complex structure induced by $J_{\alpha}$ on $\H\Lambda\backslash$ H and $\nabla^{0}$ is the ()(-)-connection on $\H\Lambda\backslash$ H.
As $\nabla$ is torsion free, so is $\nabla^{0}$, therefore, we have that $[X,Y]=0$ for any $X,Y$ vector fields on $\H\Lambda\backslash$ H induced by left-invariant vector fields on $H$. In other words, $[X,Y]=0$ for any $X,Y\in\mathcal{P}^{\nabla}$. In particular, $H$ is abelian and therefore $\H\Lambda\backslash$ H is a torus.
Since $\G\Gamma\backslash$ G is diffeomorphic to $\H\Lambda\backslash$ H and both $G$ and $H$ are nilpotent, then $G$ is isomorphic to $H$ as a Lie group [Mal]. It then follows that $G$ is abelian, and so $\G\Gamma\backslash$ G is also a torus. ∎
@@
Theorem 6.9 still holds true if instead of considering a 2-step nilmanifold we assume that $\G\Gamma\backslash$ G is a solvmanifold with $G$ a completely solvable almost abelian Lie group. Indeed, in this case the Obata connection is automatically flat, according to [AB1, Proposition 3.7], and it is complete, due to [AB1, Corollary 3.8]. For the last paragraph, we use the Saito rigidity theorem [Sai], which extends the result of Mal’cev from nilpotent to completely solvable Lie groups.
@@ 7 Finale
If $(\mathfrak{g},\{J_{\alpha}\})$ is a hypercomplex Lie algebra such that $\mathfrak{g}$ is $2$-step nilpotent, we have seen that the Obata connection is flat when at least one complex structure $J_{\alpha}$ satisfies
$J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}.$ (7.1)
We now investigate to what extent this condition can be extended to Lie algebras of higher nilpotency step. It turns out that (7.1) forces the Lie algebra to be at most $3$-step nilpotent.
Lemma 7.1.
  1. $(1)$
    Let $(\mathfrak{g},J)$ be a Lie algebra endowed with a complex structure $J$ satisfying condition (7.1). Then $\mathfrak{g}$ is at most $3$-step nilpotent and $\mathfrak{g}^{1}$ is abelian.
  2. $(2)$
    Let $(\mathfrak{g},\{J_{1},J_{2},J_{3}\})$ be a hypercomplex Lie algebra and assume that $(\alpha,\beta,\gamma)$ is a cyclic permutation of $(1,2,3)$ such that $J_{\alpha}$ and $J_{\beta}$ satisfy (7.1). Then $J_{\gamma}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$ and $\nabla_{X}Y=0$ for any $Y\in\mathfrak{g}^{1}$, where $\nabla$ is the Obata connection.
In case (1), by the integrability of $J$ and the assumption $J\mathfrak{g}^{1}\subset\mathfrak{z}$, one obtains
$X,Y],Z]=-J[[X,Y],JZ],\quad\text{for any }X,Y,Z\in\mathfrak{g},$ (7.2)
@@ 2. $(2)$
To prove (2), we may assume that $J_{1}$ and $J_{2}$ satisfy (7.1). We start by showing that $J_{3}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$, that is, $X,Y],J_{3}[U,Z]]=0$ for any $X,Y,U,Z\in\mathfrak{g}$. Indeed, using the integrability of $J_{1}$:
$[[X,Y],J_{3}[U,Z]]$ $=-J_{1}[J_{1}[X,Y],J_{3}[U,Z]]-J_{1}[[X,Y],J_{1}J_{3}[U,Z]]+[J_{1}[X,Y],J_{1}J_{3}[U,Z]]$
$=-J_{1}[J_{1}[X,Y],J_{3}[U,Z]]+J_{1}[[X,Y],J_{2}[U,Z]]-[J_{1}[X,Y],J_{2}[U,Z]]$
$=0,$
since $J_{1}\mathfrak{g}^{1}\subset\mathfrak{z}$ and $J_{2}\mathfrak{g}^{1}\subset\mathfrak{z}$. The second part of (2) follows by computing $\nabla_{X}[V,W]$ for any $X,V,W\in\mathfrak{g}$ using (2.2) and by applying the fact that both $J_{1}$ and $J_{2}$ satisfy (7.1) and (7.2). ∎
Since in the $2$-step case the corresponding Obata connection is flat (see Corollary 3.2), we will focus our attention on the $3$-step case. We observe that the hypercomplex structure must be non-abelian. Indeed, if $\{J_{\alpha}\}$ is abelian and $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for at least one $\alpha$, then
@@
Examples of 3-step nilpotent Lie algebras with hypercomplex structures have been given in [DF2, BD, AB2]. Among these, the only examples with non-abelian hypercomplex structures are those in [AB2]; however, those are Obata flat and none of them satisfies the hypothesis of Lemma 7.1 (2).
We proved in Theorem 6.1 that for hypercomplex 2-step nilmanifolds the holonomy of the Obata connection is an abelian Lie algebra, which is also a commutative associative algebra with trivial product. This theorem requires no additional assumption on the hypercomplex structure. Next, we prove a similar result without assuming 2-step nilpotency, but requiring an extra condition on the hypercomplex structure, namely, that $\{J_{1},J_{2},J_{3}\}$ satisfies the hypothesis of Lemma 7.1 (2).
Proposition 7.2.
Let $(\mathfrak{g},\{J_{1},J_{2},J_{3}\})$ be a $4n$-dimensional hypercomplex Lie algebra and assume that $(\alpha,\beta,\gamma)$ is a cyclic permutation of $(1,2,3)$ such that $J_{\alpha}$ and $J_{\beta}$ satisfy (7.1). Then:
  1. $(1)$
    the curvature tensor of the Obata connection is given by $R(X,Y)=-\operatorname{ad}_{[X,Y]}$, hence $(\mathfrak{g},\{J_{\alpha}\})$ is non-flat if and only if $\mathfrak{g}$ is $3$-step nilpotent;
  2. $(2)$
    the holonomy algebra $\mathfrak{hol}(\nabla)$ of the Obata connection $\nabla$ on $G$ is an abelian subalgebra of $\mathfrak{sl}(n,\mathbb{H})$, where $G$ is the connected and simply connected Lie group with Lie algebra $\mathfrak{g}$. Moreover, $A\cdot B=0$ for any $A,B\in\mathfrak{hol}(\nabla)$, therefore, $\mathfrak{hol}(\nabla)$ is a commutative associative algebra with trivial product.
To prove (1), we first show that $\nabla_{X}\nabla_{Y}Z=0$. We start by computing $\nabla_{Y}Z$ applying Formula (2.2). Since $\nabla_{X}$ commutes with $J_{\alpha}$ for all $\alpha$, Lemma 7.1 (2) implies that $\nabla_{X}\nabla_{Y}Z=0$. Therefore,
$R(X,Y)Z=-\nabla_{[X,Y]}Z=-\nabla_{Z}[X,Y]+[Z,[X,Y]]=-[[X,Y],Z],$
where the second equality follows from the fact that $\nabla$ is torsion-free.
In order to prove (2), set
$\mathfrak{b}:=\text{span}\{\operatorname{ad}_{[X,Y]}\mid X,Y\in\mathfrak{g}\}.$
Next, we show that $\mathfrak{b}$ is a commutative associative algebra with trivial product and that $\mathfrak{hol}(\nabla)=\mathfrak{b}$. In fact,
$\operatorname{ad}_{[X,Y]}\circ\operatorname{ad}_{[Z,W]}U=[[X,Y],[[Z,W],U]]=0,$
as $\mathfrak{g}^{1}$ is abelian (Lemma 7.1 (1)). Therefore, $\mathfrak{b}$ is a commutative associative algebra with trivial product.
Next, we show that $\mathfrak{hol}(\nabla)=\mathfrak{b}$. From Lemma 7.1 (2), we have that
$\nabla_{Z}\operatorname{ad}_{[X,Y]}=0.$
On the other hand,
$2\operatorname{ad}_{[X,Y]}\nabla_{Z}U$ $=\operatorname{ad}_{[X,Y]}J_{3}[{J_{1}Z},J_{2}U]=0,$
where the last equality follows from $J_{3}\mathfrak{g}^{1}\subset\mathfrak{z}(\mathfrak{g}^{1})$ (Lemma 7.1 (2)). By Theorem 6.3, $\mathfrak{hol}(\nabla)=\mathfrak{b}$, which implies that $\mathfrak{hol}(\nabla)$ is an abelian Lie subalgebra of $\mathfrak{sl}(n,\mathbb{H})$, which is also an associative commutative algebra with trivial product, concluding the proof. ∎
With techniques similar to those in Section 4, we can construct an example of a $3$-step nilpotent Lie algebra endowed with a hypercomplex structure $\{J_{\alpha}\}$ such that $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$ for each $\alpha$.
Example 7.3.
@@
$\begin{split}&de^{i}=0,\ i=1,\dots,4,\ \ de^{5}=e^{1,2}-e^{3,4},\\ &de^{i}=0,\ i=6,7,8,\ \ de^{9}=e^{1,3}-e^{4,2},\\ &de^{i}=0,\ i=10,11,12,\ \ de^{13}=-e^{2,5}+e^{3,9},\\ &de^{14}=e^{1,5}+e^{4,9},\ \ de^{15}=e^{4,5}-e^{1,9},\ \ de^{16}=-e^{3,5}-e^{2,9}.\end{split}$
@@
satisfying $J_{\alpha}\mathfrak{g}^{1}\subset\mathfrak{z}$, for each $\alpha=1,2,3$. Moreover, since $\mathfrak{g}$ is $3$-step, Proposition 7.2 implies that the Obata connection on the associated connected and simply connected nilpotent Lie group is not flat.